The farmhand is preparing feed for the animals. He wants to mix the contents of a bag containing 50% corn with one containing 90% corn to obtain a 30-pound bag containing 80% corn. How many pounds of the 50% bag should he use?
I don't really need help solving the equation as much as I need help so I know how to write out the equation. I have looked on many websites and in my textbook, but I cant seem to find a similar problem I can snatch the equation from.

Answer by Maths68(1474)


Let
He mixes 50% corn = x pound
He mixes 90% corn = 30-x pound
Mixture of 80% corn = 30 pound
50% corn * x pound + 90% corn * (30-x)=80% * 30
0.5*x+0.9*(30-x)=0.8 * 30
0.5x+27-0.9x=24
0.5x-0.9x=24-27
-0.4x=-3
-0.4x/-0.4=-3/-0.4
x=7.5
He should mix 7.5 pounds of 50% corn and (30-x=30-7.5=22.5) 22.5 pounds to get a mixture of 30 pound of 80% corn